OpenStudy (anonymous):

Which fraction is equivalent to 7/9?

OpenStudy (anonymous):

@salemlover352

OpenStudy (anonymous):

Are there choices? They would really help! @benator1256

OpenStudy (lol4112):

14/18 = 21/27 = 28/36 =70/90

OpenStudy (karolynxx):

28/36

OpenStudy (karolynxx):

7*4=28 9*4=36 so C is correct